OpenStudy (anonymous):

evaluate the definite integral from 0 to 1 of 2/(1 +1x^2) dx ?

OpenStudy (anonymous):

thiws question is a feather

OpenStudy (anonymous):

This one is a minor variation on a standard problem.\[\int\limits_{0}^{1}\frac{2dx}{1+x^2}=2\int\limits_{0}^{1}\frac{dx}{1+x^2}=2\arctan(1)-\arctan(0)=\frac { pi}{2}\]